# Proximal Interacting Particle Langevin Algorithms

### v2

@@ 2.3 Proximal Markov chain Monte Carlo
Intuitively, the proximity operator $x\mapsto\operatorname{prox}_{U}^{\lambda}(x)$ behaves similarly to a gradient mapping and moves points in the direction of the minimisers of $U$. In fact, when $U$ is differentiable, the proximal mapping corresponds to the implicit gradient step, as opposed to the explicit gradient step, which is known to be more stable [Parikh and Boyd, 2014]. In the limit $\lambda\to 0$ the quadratic penalty term dominates and the proximity operator coincides with the identity operator, i.e., $\operatorname{prox}_{U}^{\lambda}(x)=x$; in the limit $\lambda\to\infty$, the quadratic penalty term vanishes and the proximity mapping maps all points to the set of minimisers of $U$.
Consider a target density of the form $\pi(x)=\exp({-U(x)})/Z$ where $U:\mathbb{R}^{n}\to[0,\infty)$ is a convex lower semi-continuous function satisfying $\lim_{\|x\|\to\infty}U(x)=\infty$ and $Z$ is an unknown normalising constant. The main idea of proximal methods is to approximate the non-differentiable but log-concave target density $\pi\propto\exp(-U)$ by substituting the potential $U$ with a smooth approximation $U^{\lambda}$ whose level of smoothness is controlled by the proximal parameter $\lambda>0$ [Pereyra, 2016; Durmus et al., 2018]. The proximity map (Definition 2.1) allows us to define a family of approximations to $\pi$, indexed by $\lambda$ and referred to as Moreau-Yosida approximation. We provide below the definition of Moreau-Yosida approximations $U^{\lambda}$ and the corresponding approximations to the target density $\pi_{\lambda}$.
Definition 2.2 (Moreau-Yosida approximation [Moreau, 1965]).
@@
$U^{\lambda}(x)\coloneqq\min_{z\in\mathbb{R}^{d}}\left\{U(z)+\|z-x\|^{2}/(2\lambda)\right\}=U(\operatorname{prox}_{U}^{\lambda}(x))+\|\operatorname{prox}_{U}^{\lambda}(x)-x\|^{2}/(2\lambda).$
Consequently, we define the $\lambda$-Moreau-Yosida approximation of $\pi$ as the following density $\pi_{\lambda}(x)\propto\exp({-U^{\lambda}(x)})$.
The approximation $\pi_{\lambda}$ converges to $\pi$ as $\lambda\to 0$ (Rockafellar and Wets [2009, Theorem 1.25] and Durmus et al. [2018, Proposition 1]) and is differentiable even if $\pi$ is not, with log-gradient
$\nabla\log\pi_{\lambda}(x)=-\nabla U^{\lambda}(x)=(\operatorname{prox}_{U}^{\lambda}(x)-x)/\lambda$ (6)
(Rockafellar and Wets [2009, Example 10.32, Theorem 9.18]). Since $\pi_{\lambda}$ is now continuously differentiable, it has been suggested in Pereyra [2016]; Durmus et al. [2018] to use the discretisation of the Langevin diffusion associated with $\pi_{\lambda}$ given by
$\mathrm{d}\mathbf{X}_{\lambda,t}=\nabla\log\pi_{\lambda}(\mathbf{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},$ (7)
to approximately sample from $\pi$. Using (6) we can write the above as
$\mathrm{d}\mathbf{X}_{\lambda,t}=\lambda^{-1}(\operatorname{prox}_{U}^{\lambda}(\mathbf{X}_{\lambda,t})-\mathbf{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}.$
In this paper, we consider two classes of proximal Langevin algorithms, obtained by considering different ways to discretise (7): those obtained considering Euler–Maruyama discretisations [Pereyra, 2016; Durmus et al., 2018], and those based on splitting schemes [Klatzer et al., 2023; Habring et al., 2023; Ehrhardt et al., 2023; Durmus et al., 2019; Salim et al., 2019].
@@ 2.3.1 Proximal Langevin methods
Applying a simple Euler–Maruyama discretisation with time discretisation step $\gamma>0$ to (7), we obtain
$X_{n+1}=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}+\frac{\gamma}{\lambda}\operatorname{prox}_{U}^{\lambda}(X_{n})+\sqrt{2\gamma}\,\xi_{n+1}$ (8)
for $n\geq 0$, where $(\xi_{n})$ is a sequence of i.i.d. $d$-dimensional standard Gaussians.
Pereyra [2016] proposes the proximal unadjusted Langevin algorithm (PULA) to sample from $\pi$, which corresponds to setting $\lambda=\gamma$ in (8)
$X_{n+1}=\operatorname{prox}_{U}^{\gamma}(X_{n})+\sqrt{2\gamma}\,\xi_{n+1},$
for $n\geq 0$, where $(\xi_{n})$ are i.i.d. standard Gaussians. The computational performance of the algorithm depends strongly on the capacity to evaluate efficiently $\operatorname{prox}_{U}^{\gamma}$.
In the case where $U$ can be expressed as $U(x)=g_{1}(x)+g_{2}(x)$, with $g_{1},g_{2}$ lower bounded functions, $g_{1}\in\mathcal{C}^{1}$ convex and gradient Lipschitz, and $g_{2}$ proper, convex and lower semi-continuous, Durmus et al. [2018] propose to define $U^{\lambda}(x)=g_{1}(x)+g_{2}^{\lambda}(x)$. In that case, (7) can be rewritten as
$\mathrm{d}\mathbf{X}_{\lambda,t}=-(\nabla g_{1}(\mathbf{X}_{\lambda,t})+\nabla g_{2}^{\lambda}(\mathbf{X}_{\lambda,t}))\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},$
for $t\geq 0$. Using (6) for $g_{2}^{\lambda}$, we can write the above as
$\mathrm{d}\mathbf{X}_{\lambda,t}=-\nabla g_{1}(\mathbf{X}_{\lambda,t})+\lambda^{-1}(\operatorname{prox}_{g_{2}}^{\lambda}(\mathbf{X}_{\lambda,t})-\mathbf{X}_{\lambda,t})\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t}.$
Using a simply Euler-Maruyama discretisation with time step $\gamma>0$, the above leads to the Moreau-Yosida ULA (MYULA) algorithm
$X_{n+1}=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}-\gamma\nabla g_{1}(X_{n})+\frac{\gamma}{\lambda}\operatorname{prox}_{g_{2}}^{\lambda}(X_{n})+\sqrt{2\gamma}\,\xi_{n+1},$
for $n\geq 0$. Under some regularity conditions on $g_{1}$ and $g_{2}$, Durmus et al. [2018, Theorem 3] establish rigorous bounds on the convergence of the resulting Markov chain.
@@ 2.3.2 Proximal gradient MCMC methods
Another class of proximal sampling algorithms are based on the idea that in the case $U(x)=g_{1}(x)+g_{2}(x)$ with $g_{1}$ smooth and $g_{2}$ proper, convex and lower semi-continuous one can employ splitting schemes to modify the gradient in the ULA update (3).
Inspired by the proximal gradient algorithm (see, e.g., Parikh and Boyd [2014, Section 4.2] or Combettes and Pesquet [2011]), which provides a forward-backward splitting optimisation algorithm, Salim et al. [2019] present a sampling algorithm composed of a forward step for $g_{1}$ with the addition of a stochastic term (corresponding to one step of ULA with target $\exp(-g_{1})$) and a backward step using the proximity map of $g_{2}$
@@
where $(\xi_{n})_{n\in\mathbb{N}}$ is a sequence of i.i.d. $d$-dimensional standard Gaussians and $\gamma$ is the time discretisation step. Combining all the steps we obtain the proximal gradient Langevin algorithm (PGLA)
$X_{n+1}=\operatorname{prox}^{\lambda}_{g_{2}}\left(X_{n}-\gamma\nabla g_{1}(X_{n})+\sqrt{2\gamma}\xi_{n+1}\right).$
These algorithms were originally proposed as an alternative to MYULA to deal with cases in which $U$ is the sum of a differentiable likelihood $g_{1}$ and a compactly supported $g_{2}$, since the application of the proximity map after the addition of the stochastic term guarantees that $X_{n+1}$ remains in the support of $g_{2}$ [Salim and Richtarik, 2020].
The algorithm is further analysed in Salim and Richtarik [2020]. In addition, Ehrhardt et al. [2023] give conditions under which using an approximate proximity map does not affect numerical results and generalise existing nonasymptotic and asymptotic convergence bounds.
@@ 3 Proximal Interacting Particle Methods for MMLE
Our goal is to extend the interacting particle algorithms proposed for the MMLE problem (1) to cases where the distribution $p_{\theta}(x,y)$ is strongly log-concave in $\theta$ and $x$ jointly and may be non-differentiable. To achieve this, we build on the previously presented methodology, where we approximate the target distribution $\pi\propto\exp(-U)=\exp(-g_{1}-g_{2})$ with a Moreau-Yosida envelope $\pi_{\lambda}$ and we derive a numerical scheme based on the Euler-Maruyama discretisation or we employ a splitting scheme. In particular, inspired by interacting particle algorithms for MMLE [Akyildiz et al., 2023; Kuntz et al., 2023] and proximal Langevin methods, we introduce our methodology and three classes of proximal algorithms for the MMLE problem.
Recall that $U(\theta,x)=-\log p_{\theta}(x,y)$ as $y$ is fixed. We start by a remark which clarifies the meaning of the proximal map in the MMLE setting.
Remark 3.1.
In our scenario, $U$ is a function of $(\theta,x)$, therefore, the $\operatorname*{arg\,min}$ in the proximal map is taken over both variables, that is,
@@ 3.1 Proximal Interacting Particle Algorithms
Below we provide a set of algorithms each termed differently, that are based on discretisations of the SDE corresponding to the Moreau-Yosida approximations. We start by introducing the continuous-time interacting SDEs that will be discretised to obtain the algorithms:
@@
where $(\mathbf{B}_{t}^{i,N})_{t\geq 0}$ for $i=0,\ldots,N$ are a family of independent Brownian motions and $U^{\lambda}$ is the Moreau-Yosida approximation of $U$. Next, we will employ different discretisation strategies to obtain algorithms from this set of SDEs. We note that similar strategies will also apply to SDEs proposed in Kuntz et al. [2023], which we will come back to later.
As in the case of the interacting SDE in eq. (5), one can show that (9)–(10) converges to an SDE of the McKean–Vlasov type (MKVSDE) as $N\to\infty$. In particular, if the potential $U$ is regular enough, the MKVSDE that (9)–(10) approximates becomes arbitrarily close to that approximated by (5) if $\lambda\to 0$ (see Appendix C for a proof).
3.1.1 Proximal Interacting Particle Unadjusted Langevin algorithm
Using an Euler-Maruyama discretisation of (9)–(10) and using (6) for $\nabla U^{\lambda}$ as in Pereyra [2016], we obtain the following algorithm, that we name PIPULA (proximal interacting particle ULA)
$\theta_{n+1}^{N}$ $=\frac{1}{N}\sum_{i=1}^{N}\operatorname{prox}_{U}^{\gamma}(\theta_{n}^{N},X_{n}^{i,N})_{\theta}+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N},$ (11)
$X_{n+1}^{i,N}$ $=\operatorname{prox}_{U}^{\gamma}(\theta_{n}^{N},X_{n}^{i,N})_{x}+\sqrt{2\gamma}\;\xi_{n+1}^{i,N}.$ (12)
where $(\xi_{n}^{i,N})_{n\in\mathbb{N}}$ for $i=0,\ldots,N$ are i.i.d. standard Gaussians of the appropriate dimension. This algorithm is general and does not exploit any special properties of $U$.
3.1.2 Moreau-Yosida Interacting Particle Langevin Algorithm
Algorithm 1 Moreau-Yosida Interacting Particle Langevin Algorithm (MYIPLA)
Draw $(\theta_{0},\{X_{0}^{i,N}\}_{i=1}^{N})$ from $\pi_{\text{init}}$
for $n=0:K$ do
$\theta_{n+1}^{N}$ $=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}\theta_{n}^{N}+\frac{\gamma}{N}\sum_{i=1}^{N}\Big{(}-\nabla_{\theta}g_{1}(\theta_{n}^{N},X_{n}^{i,N})+\frac{1}{\lambda}\operatorname{prox}_{g_{2}}^{\lambda}(\theta_{n}^{N},X_{n}^{i,N})_{\theta}\Big{)}+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N}$
end for return $\theta_{K+1}^{N}$
On the other hand, if we consider $U^{\lambda}=g_{1}+g_{2}^{\lambda}$ as in Durmus et al. [2018], and substitute its gradient in the Euler-Maruyama discretisation of (9)–(10) we obtain MYIPLA (Moreau-Yosida interacting particle Langevin algorithm):
$\theta_{n+1}^{N}$ $=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}\theta_{n}^{N}+\frac{\gamma}{N}\sum_{i=1}^{N}\Big{(}-\nabla_{\theta}g_{1}(\theta_{n}^{N},X_{n}^{i,N})+\frac{1}{\lambda}\operatorname{prox}_{g_{2}}^{\lambda}(\theta_{n}^{N},X_{n}^{i,N})_{\theta}\Big{)}+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N},$ (13)
$X_{n+1}^{i,N}$ $=\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}^{i,N}-\gamma\nabla_{x}g_{1}(\theta_{n}^{N},X_{n}^{i,N})+\frac{\gamma}{\lambda}\operatorname{prox}_{g_{2}}^{\lambda}(\theta_{n}^{N},X_{n}^{i,N})_{x}+\sqrt{2\gamma}\;\xi_{n+1}^{i,N},$ (14)
where $(\xi_{n}^{i,N})_{n\in\mathbb{N}}$ for $i=0,\ldots,N$ are i.i.d. standard Gaussians and the notation $\operatorname{prox}_{g_{2}}^{\lambda}(\theta,X)_{\theta}$, $\operatorname{prox}_{g_{2}}^{\lambda}(\theta,X)_{x}$ refers to the $\theta$ and $x$ component of the proximal mapping $\operatorname{prox}_{g_{2}}^{\lambda}$, as mentioned in Remark 3.1. The algorithm is given in Algorithm 1.
3.1.3 Proximal Interacting Particle Gradient Langevin Algorithm
In the case in which $U^{\lambda}=g_{1}+g_{2}^{\lambda}$, inspired by the proximal gradient method [Salim et al., 2019; Ehrhardt et al., 2023], we employ a splitting scheme to discretise (9)–(10) and obtain PIPGLA (proximal interacting particle gradient Langevin algorithm). In this case, we perform one ULA step for both the $\theta$ and $x$ component using $\nabla g_{1}$ followed by a backward step using $\operatorname{prox}_{g_{2}}^{\lambda}$:
@@ 3.1.3 Proximal Interacting Particle Gradient Langevin Algorithm
where $(\xi_{n}^{i,N})_{n\in\mathbb{N}}$ for $i=0,\ldots,N$ are i.i.d. standard Gaussians of appropriate dimension.
Similarly to PGLA, this algorithm ensures that $X_{n+1}^{i,N}$ belongs to the support of $g_{2}$ for all $i=1,\dots,N$; if the parameter space $\Theta$ is convex, then also $\theta_{n+1}^{N}$ belongs to the support of $g_{2}$ since $\theta^{N}_{n+1}$ is a convex combination of elements of $\Theta$.
Using the relationship $\operatorname{prox}_{g_{2}}^{\lambda}(v)=v-\lambda\nabla g_{2}^{\lambda}(v)$ in (6), we can also write (15) in the following form
$\theta_{n+1}^{N}$ $=\theta_{n}^{N}-\frac{\gamma}{N}\sum_{j=1}^{N}\nabla_{\theta}g_{1}(\theta_{n}^{N},X_{n}^{j,N})-\frac{\lambda}{N}\sum_{j=1}^{N}\nabla_{\theta}g_{2}^{\lambda}(\theta_{n+1/2}^{N},X_{n+1/2}^{j,N})+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N},$
$X_{n+1}^{i,N}$ $=X_{n}^{i,N}-\gamma\nabla_{x}g_{1}(\theta_{n}^{N},X_{n}^{i,N})-\lambda\nabla_{x}g_{2}^{\lambda}(\theta_{n+1/2}^{N},X_{n+1/2}^{i,N})+\sqrt{2\gamma}\xi_{n+1}^{i,N},$
which shows that PIPGLA consists of a Langevin step with respect to $g_{1}$ and a gradient descent step with respect to $g_{2}^{\lambda}$ performed on the output of the Langevin step.
@@
Algorithm 2 Proximal Interacting Particle Gradient Langevin Algorithm (PIPGLA)
Draw $(\theta_{0},\{X_{0}^{i,N}\}_{i=1}^{N})$ from $\pi_{\text{init}}$
for $n=0:K$ do
end for return $\theta_{K+1}^{N}$
3.2 Proximal Particle Gradient Descent Methods
Our methods so far are mainly inspired from Akyildiz et al. [2023] that retain the noise in $\theta$-dimension which makes the entire system closer to a Langevin-type system. However, we can also consider removing the noise from the dynamics of $\theta$ and still provide methods using algorithms developed by Kuntz et al. [2023] and formulate theoretical results based on recent developments by Caprio et al. [2024].
If we remove the noise term in the dynamics of $\theta$ from (9), we obtain the following system of SDEs:

### v3

@@ 2.3 Proximal Methods
Intuitively, the proximity operator behaves similarly to a gradient map by moving points in the direction of the minimisers of $U$. In fact, when $U$ is differentiable, the proximal mapping corresponds to the implicit gradient step, as opposed to the explicit gradient step, which is known to be more stable [Parikh and Boyd, 2014]. Note that as $\lambda\to 0$, the proximity operator converges to the identity operator, while as $\lambda\to\infty$, the proximity operator maps all points to the set of minimisers of $U$.
The idea of proximal methods is to approximate the non-differentiable target density $\pi\propto e^{-U}$, where $U$ is a convex lower semi-continuous function, by substituting the potential $U$ with a smooth approximation $U^{\lambda}$, where the level of smoothness is controlled by the parameter $\lambda>0$. The proximity map allows us to define a family of approximations to $\pi$, indexed by $\lambda$, and referred to as Moreau-Yosida approximation [Moreau, 1965].
Definition 2 ($\lambda$-Moreau-Yosida approximation).
@@
$U^{\lambda}(x)$ $\coloneqq\min_{z\in\mathbb{R}^{d}}\left\{U(z)+\|z-x\|^{2}/(2\lambda)\right\}$
$=U(\operatorname{prox}_{U}^{\lambda}(x))+\|\operatorname{prox}_{U}^{\lambda}(x)-x\|^{2}/(2\lambda).$
Consequently, we define the $\lambda$-Moreau-Yosida approximation of $\pi$ as the following density $\pi_{\lambda}(x)\propto e^{-U^{\lambda}(x)}$.
The approximation $\pi_{\lambda}$ converges to $\pi$ as $\lambda\to 0$ [Rockafellar and Wets, 2009, Durmus et al., 2018] and is differentiable even if $\pi$ is not, with log-gradient
$\nabla\log\pi_{\lambda}(x)=-\nabla U^{\lambda}(x)=(\operatorname{prox}_{U}^{\lambda}(x)-x)/\lambda.$ (6)
Since $\pi_{\lambda}$ is continuously differentiable, proximal MCMC methods [Pereyra, 2016, Durmus et al., 2018] rely on discretisations of the Langevin diffusion associated with $\pi_{\lambda}$, given by replacing the drift term in Eq. (2) with $\nabla U^{\lambda}$, to approximately sample from $\pi$. We consider two classes of proximal Langevin algorithms, based on different discretisation schemes: Euler-Maruyama discretisations Pereyra [2016], Durmus et al. [2018], and splitting schemes Klatzer et al. [2024], Habring et al. [2024], Ehrhardt et al. [2024], Durmus et al. [2019], Salim et al. [2019].
@@ 2.3.1 Proximal Langevin methods
When $U(x)=-\log\pi(x)$ can be expressed as $U(x)=g_{1}(x)+g_{2}(x)$, with $g_{1},g_{2}$ convex lower bounded functions, $g_{1}$ differentiable and $g_{2}$ proper and lower semi-continuous, Durmus et al. [2018] consider $\pi_{\lambda}\propto e^{-U^{\lambda}}$ with $U^{\lambda}(x)=g_{1}(x)+g_{2}^{\lambda}(x)$, and the corresponding Langevin diffusion
$\mathrm{d}\mathbf{X}_{\lambda,t}=-(\nabla g_{1}(\mathbf{X}_{\lambda,t})+\nabla g_{2}^{\lambda}(\mathbf{X}_{\lambda,t}))\mathrm{d}t+\sqrt{2}\mathrm{d}\mathbf{B}_{t},\ t\geq 0.$
An Euler-Maruyama discretisation with step size $\gamma>0$ results in the Moreau-Yosida ULA (MYULA) algorithm.
@@ 2.3.2 Proximal gradient MCMC methods
Inspired by the proximal gradient algorithm (see, e.g., Parikh and Boyd [2014, Section 4.2] or Combettes and Pesquet [2011]), which is a forward-backward splitting optimisation algorithm, Salim et al. [2019] propose a sampling algorithm for the case $U^{\lambda}(x)=g_{1}(x)+g_{2}^{\lambda}(x)$ termed proximal gradient Langevin algorithm (PGLA). The method consists of a forward step in the direction of $\nabla g_{1}$ with an additional stochastic term, followed by a backward step using the proximity map of $g_{2}$:
@@
These algorithms were originally proposed as an alternative to MYULA to deal with cases in which $U$ is the sum of a differentiable likelihood $g_{1}$ and a compactly supported $g_{2}$, since the application of the proximity map after the addition of the stochastic term guarantees that $X_{n+1}$ remains in the support of $g_{2}$ [Salim and Richtarik, 2020]. In addition, Ehrhardt et al. [2024] give conditions under which using an approximate proximity map does not affect numerical results and generalise existing convergence bounds.
@@ 3 Proximal Interacting Particle Methods for MMLE
Our goal is to extend interacting particle algorithms for the MMLE problem (1) to cases where the distribution $p_{\theta}(x,y)=\pi(\theta,x)\propto e^{-U(\theta,x)}$ may be non-differentiable. To achieve this, we build on the previously presented methodology, approximating the target distribution $\pi\propto e^{-U}=e^{-g_{1}-g_{2}}$ with a Moreau-Yosida envelope $\pi_{\lambda}$ and deriving a numerical scheme using either an Euler-Maruyama discretisation or a splitting scheme. In particular, we introduce three classes of proximal algorithms.
Recall that $U(\theta,x)=-\log p_{\theta}(x,y)$ where $y$ is fixed. The proximal map in the MMLE setting is given below.
Remark 1.
In our scenario, the $\operatorname*{arg\,min}$ in the proximal map is taken over both variables $(\theta,x)$, that is,
@@ 3.1 Proximal Interacting Particle Algorithms
Our algorithms are based on different discretisation schemes of the following continuous-time interacting SDEs:
@@
where $U^{\lambda}=g_{1}+g_{2}^{\lambda}$ is the Moreau-Yosida approximation of $U$. As in the case of the interacting SDE in Eq. (5), we show that (7)–(8) converges to an SDE of the McKean–Vlasov type (MKVSDE) as $N\to\infty$. In particular, if the potential $U$ is regular enough, the MKVSDE that (7)–(8) approximates becomes arbitrarily close to that approximated by (5) as $\lambda\to 0$ (see Appendix C for a proof).
3.1.1 Moreau-Yosida Interacting Particle Langevin Algorithm (MYIPLA)
If we consider $U^{\lambda}=g_{1}+g_{2}^{\lambda}$ as in Durmus et al. [2018], and substitute its gradient in the Euler-Maruyama discretisation of (7)–(8) we derive MYIPLA (Moreau-Yosida Interacting Particle Langevin algorithm):
$\theta_{n+1}^{N}=$ $\Big{(}1-\frac{\gamma}{\lambda}\Big{)}\theta_{n}^{N}+\frac{\gamma}{N}\sum_{i=1}^{N}\Big{(}-\nabla_{\theta}g_{1}(\theta_{n}^{N},X_{n}^{i,N})$
$+\frac{1}{\lambda}\operatorname{prox}_{g_{2}}^{\lambda}(\theta_{n}^{N},X_{n}^{i,N})_{\theta}\Big{)}+\sqrt{\frac{2\gamma}{N}}\xi_{n+1}^{0,N},$ (9)
$X_{n+1}^{i,N}=$ $\Big{(}1-\frac{\gamma}{\lambda}\Big{)}X_{n}^{i,N}-\gamma\nabla_{x}g_{1}(\theta_{n}^{N},X_{n}^{i,N})$
$+\frac{\gamma}{\lambda}\operatorname{prox}_{g_{2}}^{\lambda}(\theta_{n}^{N},X_{n}^{i,N})_{x}+\sqrt{2\gamma}\;\xi_{n+1}^{i,N},$ (10)
where the notation $\operatorname{prox}_{g_{2}}^{\lambda}(\theta,X)_{\theta}$, $\operatorname{prox}_{g_{2}}^{\lambda}(\theta,X)_{x}$ refers to the $\theta$ and $x$ component of the proximal mapping $\operatorname{prox}_{g_{2}}^{\lambda}$, as mentioned in Remark 1. Setting $\gamma=\lambda$ and $g_{1}=0$ as in Pereyra [2016], we obtain a specific case of the previous algorithm, that we refer to as PIPULA (proximal interacting particle ULA).
3.1.2 Proximal Interacting Particle Gradient Langevin Algorithm (PIPGLA)
Inspired by the proximal gradient method [Salim et al., 2019, Ehrhardt et al., 2024], we employ a splitting scheme to discretise (7)–(8) and obtain PIPGLA (Proximal Interacting Particle Gradient Langevin algorithm). In this case, we perform one ULA step for both the $\theta$ and $x$ component using $\nabla g_{1}$ followed by a backward step using $\operatorname{prox}_{g_{2}}^{\lambda}$:
@@ 3.1.2 Proximal Interacting Particle Gradient Langevin Algorithm (PIPGLA)
Similarly to PGLA, this algorithm ensures that $X_{n+1}^{i,N}$ belongs to the support of $g_{2}$ for all $i$. If the parameter space $\Theta$ is convex, then also $\theta_{n+1}^{N}$ belongs to the support of $g_{2}$ since $\theta^{N}_{n+1}$ is a convex combination of elements of $\Theta$.
3.2 Proximal Particle Gradient Descent Methods (PPGD)
All the methods above incorporate a noise term in the $\theta$-dimension, making the system more akin to a Langevin-type system, as in Akyildiz et al. [2025]. However, we also explore the case where the noise term is removed from the dynamics of $\theta$, similar to Kuntz et al. [2023]. By removing the noise term from the dynamics of $\theta$ in (7), we obtain the following system of SDEs